OpenStudy (anonymous):

Tonja and Edward are participating in a jog-a-thon to raise money for charity. Tonja will raise $20, plus $2 for each lap she jogs. Edward will raise $30, plus $1.50 for each lap he jogs. The total amount of money each will raise can be calculated using the following expressions where n represents the number of laps run: Tonja: 20 + 2n Edward: 30 +1.50n Q: After how many laps will Tonja and Edward have raised the same amount of money? i don't know what to do first with this equation

OpenStudy (anonymous):

equate the amount 20 + 2n=30 +1.50n 2n-1.50n=10 0.5n=10 n=20

OpenStudy (anonymous):

what did you do with the 30? im confused at the subtract 1.50 on both sides.. how did you get 10? i subtracted 1.50 from 2 and got o.5

OpenStudy (anonymous):

30=20=10

OpenStudy (anonymous):

nevermind..i found it different way thx:)